OpenStudy (anonymous):

Linear Equations by Substitution (w/ Sol'n.) Please Kindly Solve :) x + 4y =8 2x + 8y = 3

OpenStudy (anonymous):

Since we are going about this using substitution, lets take that first equation and solve for x: \[x+4y = 8 \Leftrightarrow x = 8-4y\] Now we are going to substitute that value of x into the second equation: \[2x+8y = 3 \Leftrightarrow 2(8-4y)+8y = 3 \Leftrightarrow 16-8y +8y = 3\] but this leads to 16 = 3, which is a false statement. This shows that this system of equations has no solution.

OpenStudy (anonymous):

if anything isnt clear, feel free to ask me to clarify. :)

OpenStudy (anonymous):

if you look at the top equation and multiply thru by 2 you get 2x + 8y = 16 this is contridicting the other equation therefore unsolvable!

OpenStudy (anonymous):

No Solution

OpenStudy (anonymous):

No Solution